# Generating Physically Stable and Buildable Brick Structures from Text

### v2

@@ Abstract
{strip}[Uncaptioned image]
Figure 1: Overview of BrickGPT. (a) Our method generates physically stable interconnecting brick assembly structures from text descriptions through an end-to-end approach, showing intermediate brick-by-brick steps. (b) The generated designs are buildable both by hand and by automated robotic assembly. (c) We show example results with corresponding text prompts. Besides basic brick designs (top), our method can generate colored brick models (bottom right) and textured models (bottom left) with appearance descriptions. We highly recommend the reader to check our website for step-by-step videos.
@@ 1 Introduction
However, creating real-world objects with existing methods remains challenging. Most approaches focus on generating diverse 3D objects with high-fidelity geometry and appearance [88, 22], but these digital designs often cannot be physically realized due to two key challenges [47]. First, the objects may be difficult to assemble or fabricate using standard components. Second, the resulting structure may be physically unstable even if assembly is possible. Without proper support, parts of the design could collapse, float, or remain disconnected.
In this work, we address the challenge of generating physically realizable objects. We study this problem in the context of designing structures made of interlocking toy bricks, such as LEGO® blocks. These are widely used in entertainment, education, artistic creation, and manufacturing prototyping. Additionally, they can serve as a reproducible research benchmark, as all standard components are readily available. Due to the significant effort required to design brick structures manually, recent studies have developed automated algorithms to streamline the process and generate compelling results. However, existing approaches primarily create structures from a given 3D object [46] or focus on a single object category [14, 13].
Our goal is to develop a method for generating brick assembly structures directly from freeform text prompts while ensuring physical stability and buildability. Specifically, we aim to train a generative model that produces designs that are:
@@
  • Buildable: Compatible with standard interconnecting toy brick pieces and able to be assembled brick-by-brick by humans or robots.
In this work, we introduce BrickGPT with the key insight of repurposing autoregressive large language models, originally trained for next-token prediction, for next-brick prediction. We formulate the problem of brick structure design as an autoregressive text generation task, where the next-brick dimension and placement are specified with a simple textual format.
To ensure generated structures are both stable and buildable, we enforce physics-aware assembly constraints during both training and inference. During training, we construct a large-scale dataset of physically stable brick structures paired with captions. During autoregressive inference, we enforce feasibility with an efficient validity check and physics-aware rollback to ensure that the final tokens adhere to physics laws and assembly constraints.
Our experiments show that the generated designs are stable, diverse, and visually appealing while adhering to input text prompts. Our method outperforms pre-trained LLMs with and without in-context learning, and previous approaches based on 3D mesh generation. Finally, we explore applications such as text-driven brick texturing, as well as manual assembly and automated robotic assembly of our designs. Our dataset, code, and models are available at the project website: https://avalovelace1.github.io/BrickGPT/.
@@ 2 Related Work
Text-to-3D Generation.   Text-to-3D generation has seen remarkable progress in recent years, driven by advances in neural rendering and generative models. Dreamfusion [60] and Score Jacobian Chaining [78] pioneer zero-shot text-to-3D generation by optimizing neural radiance fields [53] with pre-trained diffusion models [64]. Subsequent work has explored alternative 3D representations [36, 3, 49, 68, 34, 43, 52] and improved loss functions [80, 86, 48, 76, 26, 45]. Rather than relying on iterative optimization, a promising alternative direction trains generative models directly on 3D asset datasets, with various backbones including diffusion models [65, 92, 33, 55, 88, 35, 21, 89, 62], large reconstruction models [22, 32, 84, 75], U-Nets [70, 37], and autoregressive models [56, 4, 69, 66, 5, 6, 19, 82].
However, these existing methods cannot be directly applied to generating brick structures because they do not account for the unique physical constraints and assembly requirements of real-world designs [47]. Our work bridges this gap by introducing a method for generating physically stable and buildable brick structures directly from text prompts.
@@
Brick Assembly and Design Generation.   Creating brick structures given a reference 3D shape has been widely studied [29]. Existing works [58, 71, 91] formulate the generation as an optimization problem guided by hand-crafted heuristic rules. Such heuristics can include ensuring that all bricks are interconnected, minimizing the number of bricks, and maximizing the number of brick orientation alternations. Wang et al. [79] translate a visual manual into step-by-step brick assembly instructions. Luo et al. [46] leverage structural stability estimation to find weak structural parts and rearrange the local brick layout to generate physically stable designs. Kim et al. [28], Liu et al. [40] formulate a planning problem to fill the target 3D model sequentially. However, these methods only generate designs given an input 3D shape, assuming a valid brick structure exists, which is difficult to verify in practice.
@@
Few works have explored learning-based techniques to generate toy brick designs. Thompson et al. [72] use a deep graph generative model in which the graph encodes brick connectivity. However, this method is limited to generating simple classes, such as walls and cuboids, using a single brick type. More recently, Ge et al. [14] use a diffusion model to predict a semantic volume, which is then translated into a high-quality micro building. Their method produces impressive results for a single category. Zhou et al. [90] and Ge et al. [13] generate compelling figurine designs given an input portrait. They use machine learning to select from a pre-made set of components that best match an input photo. Although effective for faces, extending this selection-based approach to arbitrary objects is challenging. Zhou et al. [93] formulate an optimization problem to create a brick model from an input image. While their output is a 2D brick mosaic, we focus on 3D structures in this work. Goldberg et al. [15] query a vision-language model to generate diverse 3D assembly structures. However, they use regular building blocks instead of bricks with interlocking connections, and thus the structures have limited expressiveness.
Our goal is closest to that of [31]. This work has three steps: (1) generating an image using a text-to-image model, (2) converting the image into voxels, and (3) using heuristics to create a physical brick model without considering physical constraints. In contrast, our method performs the text-to-brick task without requiring intermediate image or voxel representations.
@@
Physics-Aware Generation.   Physics-aware 3D generation can be broadly categorized into two approaches: direct constraint enforcement and learned validation. Simple physical constraints, such as collision avoidance and contact requirements, can be incorporated directly through explicit penalty terms during optimization [54, 85, 87, 77, 42, 25, 17]. More complex physical properties, such as structural stability and dynamic behavior, typically require physics simulators [12, 57, 83, 51] or data-driven physics-aware assessment models [10, 50]. To our knowledge, our paper is the first attempt to incorporate physics-aware constraints into text-based brick assembly structure generation.
@@ 3 Dataset
Training a modern autoregressive model requires a large-scale dataset. Therefore, we introduce StableText2Brick, a new large-scale dataset that contains $47{,}000+$ different toy brick assembly structures, covering $28{,}000+$ unique 3D objects from $21$ common object categories of the ShapeNetCore dataset [2]. We select categories featuring diverse and distinctive 3D objects while excluding those resembling cuboids. Each structure is paired with a group of text descriptions and a stability score, which indicates its physical stability and buildability. Below, we describe the dataset construction, an overview of which is given in Figure 2.
@@
Mesh-to-Brick.   We construct the dataset by converting 3D shapes from ShapeNetCore [2] into brick structures as shown in Figure 2(a). Given a 3D mesh, we voxelize and downsample it into a $20\times 20\times 20$ grid world to ensure a consistent scale, i.e., $H=W=D=20$. The brick layout is generated by a delete-and-rebuild algorithm similar to [46]. To improve data quality and diversity, we introduce randomness and generate multiple different structures for the same 3D object, as illustrated in Figure 2(b). This increases the chance of obtaining a stable structure and more diverse layouts. More details can be found in Appendix A. We use eight commonly available standard bricks: $1\times 1$, $1\times 2$, $1\times 4$, $1\times 6$, $1\times 8$, $2\times 2$, $2\times 4$, and $2\times 6$.
Stability Score.   We assess the physical stability of each structure, as illustrated in Figure 2(c), using the analysis method [38]. For a structure $B=[b_{1},b_{2},\dots,b_{N}]$, the stability score $S\in\mathbb{R}^{N}$ assigns each brick $b_{i}$ a value $s_{i}\in[0,1]$ that quantifies the internal stress at its connections. Higher scores $s_{i}$ indicate greater stability, while $s_{i}=0$ indicates an unstable brick that will cause structural failure. Calculating the stability score requires solving a nonlinear program to determine the forces acting on each brick to achieve a static equilibrium that prevents structural collapse, as detailed in Section 4.2. For typically-sized (i.e., $<200$ bricks) structures in Figure 2, stability analysis takes $\sim$0.35 seconds on average. A structure is stable if all bricks have stability scores greater than 0; we only include stable structures in the StableText2Brick dataset.
@@
Figure 2(e) shows several data samples in StableText2Brick. The rich variations within each category and the comprehensive text-brick pairs make it possible to train large-scale generative models. More insights on StableText2Brick are discussed in Appendix A.
@@ 4.1 Model Fine-tuning
Instruction Fine-tuning Dataset.   For each stable structure and its corresponding captions, we construct an instruction in the following format: “(user) Create a LEGO model of {caption}. (assistant) {brick-structure}.”
To simplify training and reuse LLaMA’s tokenizer, we represent brick structures in plain text. But what format should we use? The standard format LDraw [30] has two main drawbacks. First, it does not directly include brick dimensions, which are crucial for assessing the structure and validating brick placements. Second, it contains unnecessary information, such as brick orientation and scale. This information is redundant, as each axis-aligned brick has only two valid orientations.
@@ 4.2 Integrating Physical Stability
A brick structure is considered physically stable and buildable if it does not collapse when built on a baseplate. To this end, we assess physical structural stability using the stability analysis method [38]. We briefly overview this method below. Figure 4(a) illustrates all possible forces exerted on a single brick. We extend the single brick model and derive the structural force model $\mathcal{F}$, which consists of a set of candidate forces (e.g., pulling, pressing, supporting, dragging, normal, etc.), as shown in Figure 4(b). For a brick structure $B=[b_{1},b_{2},\dots,b_{N}]$, each brick $b_{i}$ has $M_{i}$ candidate forces $F_{i}^{j}\in\mathcal{F}_{i},j\in[1,M_{i}]$. A structure is stable if all bricks can reach static equilibrium, i.e.,
@@
When to apply stability analysis?   Our model generates structures sequentially, one brick at a time. A straightforward approach to ensuring physical stability is to apply stability analysis to each step and resample a brick that would cause a collapse. However, this step-by-step validation, though efficient per check, could be time-consuming due to the large number of checks required. More importantly, many structures are unstable when partially constructed, yet become stable when fully assembled. Adding a stability check after each brick generation could overly constrain the model exploration space. Instead, we propose brick-by-brick rejection sampling combined with physics-aware rollback to balance stability and diversity.
Brick-by-Brick Rejection Sampling.   To improve inference speed and avoid overly constraining the model generation, we relax our constraints during inference. First, when the model generates a brick and its position, the brick should be well-formatted (e.g., available in the inventory) and not lie outside the workspace. Second, we ensure that newly added bricks do not collide with the existing structure. Formally, for each generated brick $b_{t}$, we have $\mathcal{V}_{t}\cap\mathcal{V}_{i}=\varnothing,\forall i\in[1,t-1]$, where $\mathcal{V}_{i}$ denotes the voxels occupied by $b_{i}$. These heuristics allow us to efficiently generate well-formatted brick structures without explicitly considering complex physical stability. To integrate these heuristics, we use rejection sampling: if a brick violates the heuristics, we resample a new brick from the model. Due to the relaxed constraints, most bricks are valid, and rejection sampling does not significantly affect inference time.
@@ 1 $B\leftarrow\text{empty brick structure}$;
[…]
@@ 14                   $B\leftarrow[b_{1},\dots,b_{i-1}]$;
Table 1: Quantitative Analysis. We evaluate our method against several baselines on validity (no out-of-library, out-of-bounds, or colliding bricks), stability, CLIP-based text similarity, and DINOv2-based image similarity. Stability, CLIP, and DINO are computed over valid structures only. For LLaMA-Mesh [81], validity requires a well-formed OBJ file. Our method outperforms all baselines as well as the ablated setups on validity and stability using our proposed rejection sampling and rollback, while maintaining high text similarity.
Method % valid % stable mean brick stability min brick stability CLIP DINO
@@ 4.3 Brick Texturing and Coloring
Figure 7: Brick Texture and Color Generation. Our method can generate diverse textured (top two rows) and colored (bottom) brick structures based on the same shape while using different appearance text prompts.
@@ 5.1 Implementation Details
Fine-tuning.   Our fine-tuning dataset contains 240k distinct prompts and 47k+ distinct brick structures. We use 90% of the data for training and 10% for evaluation. For efficiency, we include samples only up to 4096 tokens in length. Training details are provided in Appendix A.
@@ 5.2 Brick Structure Generation Results
Baseline Comparisons.   As baselines, we use LLaMA-Mesh [81], LGM [70], XCube [62], and Hunyuan3D-2 [89] to generate a mesh from each prompt, then convert the meshes to brick structures with our delete-and-rebuild algorithm. Additionally, we compare our method with pre-trained LLaMA models evaluated in both a zero-shot and few-shot manner. For few-shot evaluation, we provide the model with 5 examples of stable brick structures and their captions.
For each method, we compute the proportion of stable and valid structures among the generated designs. Additionally, for each valid structure, we compute its mean and minimum brick stability scores. To evaluate prompt alignment, we compute the CLIP score [61] between a rendered image of each valid structure and the text “A LEGO model of {prompt}”. We also calculate the alignment between rendered images of the generated structure and the ground-truth structure for the same prompt, as measured by the cosine similarity between DINOv2 features [59]. As shown in Table 1, our method outperforms all baselines in these metrics. Figure 5 shows that our method generates brick structures of higher quality than the baselines.
Ablation Study.   We demonstrate the importance of rejection sampling and physics-aware rollback. As seen in Figure 6, rejection sampling eliminates invalid bricks, such as those with collisions, while rollback helps to ensure the final assembly structure is physically stable. The quantitative results in Table 1 show that our full method generates a higher proportion of valid and stable brick structures, while closely following the text prompts.

### v3

@@ 1 Introduction
However, creating real-world objects with existing methods remains challenging. Most approaches focus on generating diverse 3D objects with high-fidelity geometry and appearance [88, 22], but these digital designs often cannot be physically realized due to two key challenges [47]. First, the objects may be difficult to assemble using standard components. Second, the resulting structure may be physically unstable even if assembly is possible. Without proper support, parts of the design could collapse, float, or be disconnected.
In this work, we address the challenge of generating physically realizable objects. We study this problem in the context of designing structures made of interlocking toy bricks, such as LEGO® blocks. These are widely used in education, artistic creation, and manufacturing prototyping. Additionally, they can serve as a reproducible benchmark, as all standard components are readily available. Due to the significant effort required to design brick structures manually, recent studies have developed automated algorithms to streamline the process and generate compelling results. However, existing approaches primarily create structures from a given 3D object [46] or focus on a single object category [14, 13].
Our goal is to generate brick assembly structures directly from freeform text prompts while ensuring physical stability and buildability. Specifically, we aim to train a generative model that produces designs that are:
@@
  • Buildable: Compatible with standard toy brick pieces and able to be assembled brick-by-brick by humans or robots.
In this work, we introduce BrickGPT with the key insight of repurposing autoregressive large language models, originally trained for next-token prediction, for next-brick prediction. We formulate the problem of brick structure design as an autoregressive text generation task, where the next-brick dimension and placement are specified with a simple textual format. To ensure generated structures are both stable and buildable, we enforce physics-aware assembly constraints during both training and inference. During training, we construct a large-scale dataset of physically stable brick structures paired with captions. During autoregressive inference, we enforce feasibility with an efficient validity check and physics-aware rollback to ensure that the final tokens adhere to physics laws and assembly constraints.
Our experiments show that the generated designs are stable, diverse, and visually appealing while adhering to input prompts. Our method outperforms pre-trained LLMs with or without in-context learning, and previous approaches based on 3D mesh generation. Finally, we explore applications such as text-driven brick texturing, as well as manual assembly and automated robotic assembly of our designs. Our dataset, code, and models are available at the project website: https://avalovelace1.github.io/BrickGPT/.
@@ 2 Related Work
Text-to-3D Generation. Text-to-3D generation has seen remarkable progress recently, driven by advances in neural rendering and generative models. Dreamfusion [60] and Score Jacobian Chaining [78] pioneer zero-shot text-to-3D generation by optimizing neural radiance fields [53] with pre-trained diffusion models [64]. Subsequent work has explored alternative 3D representations [36, 3, 49, 68, 34, 43, 52] and improved loss functions [80, 86, 48, 76, 26, 45]. Rather than relying on iterative optimization, a promising alternative direction trains generative models directly on 3D asset datasets, with various backbones including diffusion models [65, 92, 33, 55, 88, 35, 21, 89, 62], large reconstruction models [22, 32, 84, 75], U-Nets [70, 37], and autoregressive models [56, 4, 69, 66, 5, 6, 19, 82]. However, these existing methods cannot be directly applied to generating brick structures as they do not account for the unique physical and assembly constraints of real-world designs [47]. We bridge this gap by introducing a method for generating physically stable and buildable brick structures directly from text prompts.
@@
Brick Assembly and Design Generation. Creating brick structures given a reference 3D shape has been widely studied [29]. Existing works [58, 71, 91] formulate the generation as an optimization problem guided by hand-crafted heuristics, such as ensuring that all bricks are interconnected and minimizing the number of bricks. Wang et al. [79] translate a visual manual into step-by-step brick assembly instructions. Luo et al. [46] use stability analysis to find weak structural parts and rearrange the local brick layout to generate physically stable designs. Kim et al. [28], Liu et al. [40] formulate a planning problem to fill the target 3D model sequentially. However, these methods only generate designs given an input 3D shape, assuming a valid brick structure exists, which is difficult to verify in practice.
@@
Few works have explored learning-based techniques to generate toy brick designs. Thompson et al. [72] use a deep graph generative model in which the graph encodes brick connectivity. However, this method is limited to generating simple classes using a single brick type. More recently, Ge et al. [14] use a diffusion model to predict a semantic volume, which is then translated into a high-quality micro building. Their method produces impressive results for a single category. Zhou et al. [90] and Ge et al. [13] generate compelling figurine designs given an input portrait, selecting components from a pre-made set that best match an input photo. Although effective for faces, extending this selection-based approach to arbitrary objects is challenging. Zhou et al. [93] formulate an optimization problem to create a brick model from an input image. While their output is a 2D brick mosaic, we focus on 3D structures in this work. Goldberg et al. [15] query a vision-language model to generate diverse 3D assembly structures. However, they use regular building blocks instead of bricks with interlocking connections, and thus the structures have limited expressiveness. Lennon et al. [31] voxelize an input image and use heuristics to convert the voxels into a brick model without considering physical constraints. In contrast, our method performs the text-to-brick task without requiring intermediate voxel representations.
@@
Physics-Aware Generation. Physics-aware 3D generation can be broadly categorized into two approaches: direct constraint enforcement and learned validation. Simple physical constraints, such as collision avoidance and contact requirements, can be incorporated directly through explicit penalty terms during optimization [54, 85, 87, 77, 42, 25, 17]. More complex physical properties, such as structural stability and dynamic behavior, typically require physics simulators [12, 57, 83, 51, 15] or data-driven physics-aware assessment models [10, 50]. To our knowledge, our paper is the first attempt to incorporate physics-aware constraints into text-based brick assembly structure generation.
@@ 3 Dataset
Training a modern autoregressive model requires a large-scale dataset. Therefore, we introduce StableText2Brick, a new dataset that contains 47,000+47{,}000+ different toy brick assembly structures, covering 28,000+28{,}000+ unique 3D objects from $21$ common object categories of the ShapeNetCore dataset [2]. We select categories featuring diverse and distinctive 3D objects while excluding those resembling cuboids. Each structure is paired with a group of text descriptions and a stability score, which indicates its physical stability and buildability. Below, we describe the dataset construction, an overview of which is given in Figure 2.
@@
Shape-to-Brick. We construct the dataset by converting 3D shapes from ShapeNetCore [2] into brick structures as shown in Figure 2(a). Given a 3D mesh, we voxelize and downsample it into a $20\times 20\times 20$ grid world to ensure a consistent scale, i.e., $H=W=D=20$. The brick layout is generated by a delete-and-rebuild algorithm similar to [46]. To improve data quality and diversity, we introduce randomness and generate multiple different structures for the same 3D object, as illustrated in Figure 2(b). This increases the chance of obtaining a stable structure and more diverse layouts. More details can be found in Appendix A.
Stability Score. We assess the physical stability of each structure, as illustrated in Figure 2(c), using the analysis method [38]. For a structure $B=[b_{1},b_{2},\dots,b_{N}]$, the stability score $S\in\mathbb{R}^{N}$ assigns each brick $b_{i}$ a value $s_{i}\in[0,1]$ that quantifies the internal stress at its connections. Higher scores $s_{i}$ indicate greater stability, while $s_{i}=0$ indicates an unstable brick that will cause structural failure. Calculating the stability score requires solving a nonlinear program to determine the forces acting on each brick to achieve a static equilibrium that prevents structural collapse, as detailed in Section 4.2. For typically-sized (i.e., <200<200 bricks) structures in Figure 2, stability analysis takes $\sim$0.35 seconds on average. A structure is stable if all bricks have stability scores greater than 0; we only include stable structures in the StableText2Brick dataset.
@@
Figure 2(e) shows several data samples from StableText2Brick. The rich variations within each category and the comprehensive text-brick pairs make it possible to train large-scale generative models. More insights on StableText2Brick are discussed in Appendix A.
@@ 4.1 Model Fine-tuning
Instruction Fine-tuning Dataset. For each stable structure and its corresponding captions, we construct an instruction in the following format: “(user) Create a LEGO model of {caption}. (assistant) {brick-structure}.” To simplify training and reuse LLaMA’s tokenizer, we represent brick structures in plain text. But what format should we use? The standard format LDraw [30] has two main drawbacks. First, it does not directly include brick dimensions, which are crucial for assessing the structure and validating brick placements. Second, it includes unnecessary information such as arbitrary brick orientation and scale, which are redundant since all bricks are axis-aligned in our setting.
@@ 4.2 Integrating Physical Stability
A brick structure is considered physically stable and buildable if it does not collapse when built on a baseplate. To this end, we assess physical structural stability using the stability analysis method [38]. We briefly overview this method below. Figure 4(a) illustrates all possible forces exerted on a single brick. We extend the single brick model and derive the structural force model $\mathcal{F}$, which consists of a set of candidate forces as shown in Figure 4(b). For a brick structure $B=[b_{1},b_{2},\dots,b_{N}]$, each brick $b_{i}$ has $M_{i}$ candidate forces $F_{i}^{j}\in\mathcal{F}_{i},j\in[1,M_{i}]$. A structure is stable if all bricks can reach static equilibrium, i.e.,
@@
When to apply stability analysis? A straightforward approach to ensuring physical stability is to apply stability analysis after each generated brick and resample a brick that would cause a collapse. However, this step-by-step validation could be time-consuming. More importantly, many structures are unstable during construction yet become stable when fully assembled; adding a stability check after each brick generation could overly constrain the model exploration space. Instead, we propose brick-by-brick rejection sampling combined with physics-aware rollback to balance stability and diversity.
Brick-by-Brick Rejection Sampling. To improve inference speed and avoid overly constraining generation, we relax our constraints during inference. First, when the model generates a brick and its position, the brick should be well-formatted (e.g., available in the inventory) and lie within the world space. Second, the brick should not collide with the existing structure. Formally, for each generated brick $b_{t}$, we have $\mathcal{V}_{t}\cap\mathcal{V}_{i}=\varnothing,\forall i\in[1,t-1]$, where $\mathcal{V}_{i}$ denotes the voxels occupied by $b_{i}$. These heuristics allow us to efficiently generate well-formatted structures without explicitly considering complex physical stability. To integrate these heuristics, we use rejection sampling: if a brick violates the rules, we resample a new brick from the model. Due to the relaxed constraints, most bricks are valid, and rejection sampling does not significantly affect inference time.
@@ 1 $B\leftarrow\text{empty brick structure}$;
3  for $k=1,\ldots,\texttt{max_rejections}$ do
4     $\texttt{context}\leftarrow T\oplus B.\texttt{to_text_format}()$;
5     $b\leftarrow\theta.\texttt{predict_tokens}(\texttt{context})$ (Eqn. 1);
@@ 14        $B\leftarrow[b_{1},\dots,b_{i-1}]$;
Table 1: Quantitative Results. We evaluate our method against several baselines on validity (no out-of-library, out-of-bounds, or colliding bricks), stability, CLIP-based text similarity, and DINOv2-based image similarity. Stability, CLIP, and DINO are computed over valid structures only. For LLaMA-Mesh [81], validity requires a well-formed OBJ file. Results marked “+ our stability analysis” are augmented by generating multiple structures and choosing the first stable one found (if any). Our method outperforms all baselines as well as the ablated setups on validity and stability using our proposed rejection sampling and rollback, while maintaining high text similarity.
Method % valid % stable mean stability min stability CLIP DINO
@@
LLaMA-Mesh [81] + our stability analysis 94.8% 58.0% 0.896 0.564 0.317 0.851
LGM [70] + our stability analysis 100% 32.5% 0.941 0.285 0.300 0.851
XCube [62] + our stability analysis 100% 83.6% 0.963 0.754 0.322 0.859
Hunyuan3D-2 [89] + our stability analysis 100% 88.4% 0.976 0.813 0.324 0.868
@@ 5.1 Implementation Details
Fine-tuning. Our fine-tuning dataset contains 240k distinct prompts and 47k+ distinct brick structures. We withhold 10% of the data for evaluation. For efficiency, we include samples only up to 4096 tokens long. Training details are provided in Appendix A.
@@ 5.2 Brick Structure Generation Results
Figure 7: Brick Texture and Color Generation. Our method can generate diverse textured (top two rows) and colored (bottom) structures based on the same shape with different appearance prompts.
Baseline Comparisons. As baselines, we use LLaMA-Mesh [81], LGM [70], XCube [62], and Hunyuan3D-2 [89] to generate a mesh from each prompt, then convert the meshes to brick structures with our delete-and-rebuild algorithm. To increase the chance of producing a stable structure, we generate multiple different structures for the same output mesh (10 for LLaMA-Mesh and LGM; 100 for XCube and Hunyuan3D) and accept the first stable one found, if any. Additionally, we compare our method with pre-trained LLaMA models evaluated in both a zero-shot and few-shot manner. For few-shot evaluation, we provide the model with 5 examples of stable brick structures and their captions.
For each method, we compute the proportion of stable and valid structures among the generated designs. Additionally, for each valid structure, we compute its mean and minimum brick stability scores. To evaluate prompt alignment, we compute the CLIP score [61] between a rendered image of each valid structure and the text “A LEGO model of {prompt}”. We also calculate the alignment between rendered images of the generated structure and the ground-truth structure for the same prompt, as measured by the cosine similarity between DINOv2 features [59]. As shown in Table 5, our method outperforms all baselines in these metrics.
Ablation Study. We demonstrate the importance of rejection sampling and physics-aware rollback. As seen in Figure 6, rejection sampling eliminates invalid bricks, such as those with collisions, while rollback helps to ensure the final assembly structure is stable. The quantitative results in Table 5 show that our full method generates a higher proportion of valid and stable brick structures, while closely following the text prompts.